# CliquePH: Higher-Order Information for Graph Neural Networks through Persistent Homology on Clique Graphs

### v1

@@ 2 Preliminaries
Persistent homology [10, 11], a technique for capturing topological features of data, constitutes the foundation of our proposed method. In contrast to other graph learning techniques, persistent homology permits us to quantify topological features at different scales. The crucial element enabling multi-scale calculations is a filtration, i.e., a consistent ordering of the vertices and edges of a graph, defined using a function $f\colon G\to\mathbb{R}$. Filtrations can be either obtained from static descriptor functions that measure certain aspects of a graph, such as its number of connected components or its curvature [12], but recent work also demonstrates that it is possible to learn filtrations in an end-to-end fashion [13, 9]. Regardless of the origin of a filtration, the crucial insight is that, since a graph is a discrete object, its topology can only change at a finite number of critical thresholds $t_{1},\dotsc,t_{k}$. Thus, any function $f\colon G\to\mathbb{R}$ that gives rise to a filtration enables us to turn a graph into a sequence of nested subgraphs $G_{1}\subseteq G_{2}\subseteq\dotsc\subseteq G_{k}=G$, where each $G_{i}$ is defined based on the critical points of $f$ via $G_{i}:=\{x\in V\cup E\mid f(x)\leq t_{i}\}$. Intuitively, each subgraph consists of all elements of the graph $G$ whose function value is less than or equal to the critical threshold. Alongside this sequence of subgraphs, we now calculate the homology groups. These groups capture the topological features—connected components, cycles, voids, and higher-order features—of each subgraph. Persistent homology represents all topological features arising from the filtration in a set of persistence diagrams. Each persistence diagram consists of a multiset of intervals of the form $(c_{i},d_{i})$ with $c_{i},d_{i}\in\mathbb{R}\cup\{\infty\}$, with $c_{i}$ denoting the “creation time” (in terms of the critical thresholds ) and $d_{i}$ denoting the “destruction time.” Features with an infinite destruction time are called essential; these denote topological features that appear at all scales. Prior work in leveraging topological features from graphs ignored clique information due to computational issues: while it is possible to work with higher-order topological features arising from cliques, a naive application of the persistent homology scales with $\mathcal{O}(n^{d})$, where $n$ denotes the number of vertices in a graph, and $d$ denotes the order of cliques (we refer to this as exact higher order persistent homology). Our proposed method instead lifts cliques into individual skeleton graphs, and then performs the efficient persistent homology up to dimension 1. In Table 1 we show a time comparison for one epoch of training, of our proposed method CliquePH with a lifting of the original graph up to 3-vertex clique graphs, and a modified version making use of exact persistent homology up to dimension 2 (dimension 2 corresponds to triangles, i.e., 3-vertex cliques – hence the comparison). The exact persistent homology computations take about 190 times more than our method CliquePH [^1] The output of CliquePH is not the same as performing exact higher-order persistent homology; Table 1 is meant to show the impracticality of performing the latter..
@@ 3 Our Method
Our proposed method aims at introducing information from persistent homology on higher-order structures into GNNs. While our method is theoretically different from performing persistent homology up to higher dimensions, we show it brings substantial performance improvements. Our method can be decomposed into four components (as shown in Figure 1), which we present below.
@@ 3.1 Graph Lifting
Starting from the original graph $G$, we perform a “lifting“ operation to extract its clique graphs:
@@
In practice, we use $r=5$ in our experiments as available datasets tend to have a very small number of cliques larger than that. When dealing with real-world graphs, several works have proposed methods for enumerating all cliques in a graph that in practice have a fast runtime [15, 16, 17, 18]. In all datasets considered in our experiments, we have been able to perform this pre-processing step for each dataset, up to $r=5$, within 1 hour. For instance, complete preprocessing time on MNIST dataset for up to 3-cliques, up to 4-cliques, and up to 5-cliques takes 36.4 min, 58.1 min, and 60.7 min consecutively.
@@ 4 Theoretical Motivation
As CliquePH makes use of first-order persistent homology information, it is straightforward to see from Theorem 4.1 that a GNN empowered with CliquePH is strictly more powerful than the 1-WL algorithm. While we cannot use Theorem 4.2 to theoretically prove that a version of CliquePH with cliques up to order $k$ leads to the expressiveness of $k$-WL (as CliquePH does not perform exact higher-order persistent homology), we provide below empirical evidence of the practical increase in expressiveness provided by CliquePH.
@@
Table 3: Number of strongly regular graphs distinguishable by untrained models. All models have the same number of layers and dimensionality. CliquePH can enhance the ability of networks to distinguish “hard” non-isomorphic graphs.
@@ 5 Experiments
To demonstrate the benefits provided by our method in practical scenarios, we follow prior literature [9] and apply our method to three GNNs among the most used by practitioners: GCN [23], GIN [6], and GAT [24]. In more detail, we report results for each GNN by itself, for the GNN with the addition of TOGL [9], and for the GNN with the addition of our method CliquePH. The comparison with TOGL is particularly useful, as TOGL is a topological method that can be used on top of any GNN (like CliquePH), and it makes use of information coming from persistent homology up to dimension one. This then allows us to analyze the benefit provided by the persistent homology information obtained from the higher-order clique graphs that is present in our method. Finally, we provide an ablation study to observe the effects of the position of the CliquePH layer in the model architecture.
@@
In order to ensure a fair performance comparison, we take the optimal architectures and hyperparameters for the GNNs and for TOGL from prior work [9]. For CliquePH we use the same architecture for the base GNN of TOGL, and we tune only the learning rate and the maximum dimension of the lifted clique graphs (between 3, 4, and 5) using a random search approach (using values in the interval: $(10^{-6},10^{-2})$) by comparing the loss on the validation set. Furthermore, the reported results are obtained by averaging over ten independent runs with varying random seeds. The code to replicate our experiments will be publicly released upon acceptance. In Tables 2 and 4, the notation “GCN-4” indicates a GCN architecture with 4 message passing layers, while “GCN-3-CliquePH-1” indicates a GCN architecture with 3 message passing layers, and 1 CliquePH layer. We use the analogous notation for the other GNNs, and for networks with TOGL. All experimental results were conducted on a single NVIDIA A100 GPU, and 20 allocated AMD EPYC 7742 64-Core CPUs, with the Adam optimizer, 4 workers, and a batch size of 32 for small datasets and 128 for large datasets. We use standard graph classification benchmark datasets: DD [25], ENZYMES [26], MNIST [27], PROTEINS [28], IMDB-B [29], Reddit-5k [30], and OGBG-Molhiv [31]. We perform two types of experiments, the first, which we refer to as structure-based, does not consider the information coming from node features; the second instead follows the standard procedure, and makes use of the available node attributes. To ensure the same training, evaluation, and data setup, we use the code provided by the authors of TOGL [9], and we report results for TOGL and base models from their paper.
@@
Following prior work [9], to perform structure-based experiments, we assign random node labels and input features. This approach allows us to assess the practical effectiveness of our method in comparison to classical message-passing GNNs for extracting underlying structural and topological information. Table 2 shows the results for the structure-based graph classification task. Firstly, we observe that CliquePH and TOGL significantly improve the performance of standard GNNs. CliquePh in particular leads to the highest average improvement across datasets (last column on the right), with an improvement of $6.9\%$ with respect to the base GCN model, $2.3\%$ with respect to the base GIN, and $10.7\%$ with respect to the base GAT. Looking at individual datasets, we then notice that a CliquePH model is the best performing model on 6 out of 7 datasets. Furthermore, it can be seen that in most cases, the results for CliquePH present a lower variance across runs, showing an increased stability with respect to TOGL.
@@
We investigate the effects of the position of the CliquePH layer inside a GNN architecture, and the maximum dimension $r$ of the lifted clique graphs. We provide below an overview of the results and we refer to Appendix C for the full results. We start by considering a 4-layer GNN, with 3 message passing layers and 1 CliquePH layer, and we change the position of the latter. We consider the structure-based scenario to exclude the influence of node features. Results are shown in Figure 2 (Left) for the DD dataset (other datasets are in the Appendix). We notice that the effects are highly dataset dependent: for DD using CliquePH as the first layer is more effective, while, e.g., for Enzymes the best performance is when the layer is 3rd, similar to PROTEINS. We then consider a GNN with a number of layers that vary from 1 to 5 and with CliquePH as the last layer. Results are shown in Figure 2 (Center). Again we notice that the performance is highly dependent on the dataset, as for DD there is little change in performance for different numbers of layers, while, e.g., for ENZYMES it is important to have 2 to 4 layers prior to CliquePH. On IMDB-B we notice how the addition of CliquePH significantly increases performance even for a 1-layer architecture. Finally, we notice how in all cases the addition of CliquePH always leads to performance improvements. Finally, we apply CliquePH with a maximum dimension of the lifted clique graphs varying from 3 to 5 on the DD dataset. Results are shown in Figure 2 (Right). We observe that tuning this parameter is important to avoid overfitting. These results suggest that tuning the position of CliquePH, and the value of $r$, is important to obtain the best results, and that CliquePH always leads to performance improvements when no node features are available.
@@ 6 Related Work
Graph neural networks and topological data analysis have become very popular and prolific research fields. In this section, we review works at the intersection of graph machine learning and topology, and we refer the interested reader to comprehensive surveys for the latest advancements in graph neural networks [1, 2] and topological data analysis [32, 33, 34]. In the context of graph neural networks augmented with topological information, Bouritsas et al. [7] show that adding information about the number of higher-order structures that a node belongs to can improve the theoretical expressivity and practical performance of a model. Similarly, Immonen et al. [35] use a topological descriptor based on filtrations as additional node features for GNNs. Zhao et al. [36] use persistent homology to weight messages in GNNs in the message passing procedure. He et al. [37] use sheaf theory to produce node position encodings that improve the performance of graph neural networks on several tasks. Song et al. [38] propose a feature augmentation method by obtaining topology embedding of nodes through Node2vec. This introduces topological structure information into an end-to-end model to improve node representation learning. In addition, several works discuss message-passing on topological spaces. Bodnar et al. [39] introduce a model that performs message passing over simplicial complexes. Bodnar et al. [40] successively extended this model to cellular complexes. Later work [41] has also introduced models that operate on cellular sheaves. While these works are related to CliquePH, there are some important differences. Firstly, these methods require a specific message-passing model, while our method CliquePH can be applied in conjunction with any existing GNN. Secondly, CliquePH makes use of Persistent Homology, which provides access to information (e.g., the exact number of cliques and cycles in the original graph and in each lifted graph) that cannot be guaranteed to be accessible to methods that perform message passing on topological structures. Finally, several works also focus on the explicit connections between topological data analysis and graph learning. O’Bray et al. [42] use graph filtrations to compute efficient graph representations that can outperform those obtained from GNNs. Southern et al. [43] leverage topological tools to robustly evaluate generative graph models. Coupette et al. [44] introduces the concept of curvature for hypergraphs and show that they can be used to effectively perform clustering. Carriere et al. [45] use extended persistent diagrams to produce graph embeddings. Srambical and Rieck [46] use graph filtrations to classify dynamic graphs. Ye et al. [47] introduces the use of extended persistent homology, a variant of persistent homology, into graph-related tasks.
@@ Appendix C Full Ablation results
Appendix D Node Classification Experiments

### v2

@@ 2 Preliminaries
Persistent homology [10, 11], a technique for capturing topological features of data, constitutes the foundation of our proposed method. In contrast to other graph learning techniques, persistent homology permits us to quantify topological features at different scales. The crucial element enabling multi-scale calculations is a filtration, i.e., a consistent ordering of the vertices and edges of a graph, defined using a function $f\colon G\to\mathbb{R}$. Filtrations can be either obtained from static descriptor functions that measure certain aspects of a graph, such as its number of connected components or its curvature [12], but recent work also demonstrates that it is possible to learn filtrations in an end-to-end fashion [13, 9]. Regardless of the origin of a filtration, the crucial insight is that, since a graph is a discrete object, its topology can only change at a finite number of critical thresholds $t_{1},\dotsc,t_{k}$. Thus, any function $f\colon G\to\mathbb{R}$ that gives rise to a filtration enables us to turn a graph into a sequence of nested subgraphs $G_{1}\subseteq G_{2}\subseteq\dotsc\subseteq G_{k}=G$, where each $G_{i}$ is defined based on the critical points of $f$ via $G_{i}:=\{x\in V\cup E\mid f(x)\leq t_{i}\}$. Intuitively, each subgraph consists of all elements of the graph $G$ whose function value is less than or equal to the critical threshold. Alongside this sequence of subgraphs, we now calculate the homology groups. These groups capture the topological features—connected components, cycles, voids, and higher-order features—of each subgraph. Persistent homology represents all topological features arising from the filtration in a set of persistence diagrams. Each persistence diagram consists of a multiset of intervals of the form $(c_{i},d_{i})$ with $c_{i},d_{i}\in\mathbb{R}\cup\{\infty\}$, with $c_{i}$ denoting the “creation time” (in terms of the critical thresholds ) and $d_{i}$ denoting the “destruction time.” Features with an infinite destruction time are called essential; these denote topological features that appear at all scales. Prior work in leveraging topological features from graphs ignored clique information due to computational issues: while it is possible to work with higher-order topological features arising from cliques, a naive application of the persistent homology scales with $\mathcal{O}(n^{d})$, where $n$ denotes the number of vertices in a graph, and $d$ denotes the order of cliques [^1] This complexity is due to the number of possible simplices of dimension $d$. (we refer to this as exact higher order persistent homology). Our proposed method instead lifts cliques into individual skeleton graphs, and then performs the efficient persistent homology up to dimension 1. In Table 1 we show a time comparison for one epoch of training, of our proposed method CliquePH with a lifting of the original graph up to 3-vertex clique graphs, and a modified version making use of exact persistent homology up to dimension 2 (dimension 2 corresponds to triangles, i.e., 3-vertex cliques – hence the comparison). The exact persistent homology computations take about 190 times more than our method CliquePH [^2] The output of CliquePH is not the same as performing exact higher-order persistent homology; Table 1 is meant to show the impracticality of performing the latter..
@@ 3 Our Method
Our proposed method aims at introducing information from persistent homology on higher-order structures into GNNs. Our method can be decomposed into four components (as shown in Figure 1), which we present below.
@@ 3.1 Graph Lifting
Starting from the original graph $G$, we perform a “lifting” operation to extract its clique graphs:
@@
The motivation behind the use of clique graphs is threefold: (1) several works have proposed methods for efficiently enumerating all cliques in large graphs [15, 16, 17, 18], (2) cliques are heavily studied in graph theory and have a vast amount of applications (e.g., see [19]), and (3) cliques have already been used successfully to improve performance on GNNs [7].
In practice, we use $r=5$ in our experiments as available datasets tend to have a very small number of cliques larger than that. In all considered datasets, we have been able to perform this pre-processing step up to $r=5$, within 1 hour. For instance, complete preprocessing time on MNIST dataset for up to 3-cliques, up to 4-cliques, and up to 5-cliques takes 36.4 minutes, 58.1 minutes, and 60.7 minutes.
@@ 4 Theoretical Motivation
As CliquePH makes use of first-order persistent homology information, it is straightforward to see from Theorem 4.1 that a GNN empowered with CliquePH is strictly more powerful than the 1-WL algorithm. While we cannot use Theorem 4.2 to theoretically prove that a version of CliquePH with cliques up to order $k$ leads to the expressiveness of $k$-WL (as CliquePH does not perform exact higher-order persistent homology), we provide below some empirical evidence.
@@
Table 3: Number of strongly regular graphs distinguishable by untrained models. CliquePH can enhance the ability of networks to distinguish “hard” non-isomorphic graphs.
@@ 5 Experiments
To demonstrate the benefits provided by our method in practical scenarios, we follow prior literature [9] and apply our method to three GNNs among the most used by practitioners: GCN [24], GIN [6], and GAT [25]. We report results for each GNN by itself, for the GNN with the addition of TOGL [9], and for the GNN with the addition of our method CliquePH. The comparison with TOGL is particularly useful, as TOGL is a topological method that can be used on top of any GNN (like CliquePH), and it makes use of information coming from persistent homology up to dimension one. This then allows us to analyze the benefit provided by the persistent homology information obtained from the higher-order clique graphs that is present in our method. Finally, we provide an ablation study to observe the effects of the position of the CliquePH layer in the model architecture.
@@
In order to ensure a fair performance comparison, we take the optimal architectures and hyperparameters for the GNNs and for TOGL from prior work [9]. For CliquePH we use the same architecture for the base GNN of TOGL, and we tune only the learning rate and the maximum dimension of the lifted clique graphs (between 3, 4, and 5) using a random search approach (using values in the interval: $(10^{-6},10^{-2})$) by comparing the loss on the validation set. Furthermore, the reported results are obtained by averaging over ten independent runs with varying random seeds. The code to replicate our experiments will be publicly released upon acceptance. We provide statistics about the used models and datasets in Appendix F. In Tables 2 and 4, the notation “GCN-4” indicates a GCN architecture with 4 message passing layers, while “GCN-3-CliquePH-1” indicates a GCN architecture with 3 message passing layers, and 1 CliquePH layer. We use the analogous notation for the other GNNs, and for networks with TOGL. All experimental results were conducted on a single NVIDIA A100 GPU, and 20 allocated AMD EPYC 7742 64-Core CPUs, with the Adam optimizer, 4 workers, and a batch size of 32 for small datasets and 128 for large datasets. We use standard graph classification benchmark datasets: DD [26], ENZYMES [27], MNIST [28], PROTEINS [29], IMDB-B [30], Reddit-5k [31], and OGBG-Molhiv [32]. To ensure the same training, evaluation, and data setup, we use the code provided by the authors of TOGL [9], and we report results for TOGL and base models from their paper.
@@
Following prior work [9], we perform “structure-based” graph classification experiments, in which we assign random node features. This approach removes the information present in the features, and allows us to assess the models’ effectiveness when relying only on structural information. Results are shown in Table 2. Firstly, we observe that CliquePH and TOGL significantly improve the performance of standard GNNs. CliquePH in particular leads to the highest average improvement across datasets (right-most column), with an improvement of $6.9\%$ with respect to the base GCN model, $2.3\%$ for the base GIN, and $10.7\%$ for the base GAT. Looking at individual datasets, we then notice that a CliquePH model is the best performing model on 6 out of 7 datasets. Furthermore, it can be seen that in most cases, the results for CliquePH present a lower variance across runs, showing an increased stability with respect to TOGL.
@@
We investigate the effects of the position of the CliquePH layer inside a GNN architecture, and the maximum dimension $r$ of the lifted clique graphs. We provide below an overview of the results and we refer to Appendix C for the full results. Furthermore, in Appendix D we compare against a version CliquePH with static (non-learnable) embedding functions for the persistence diagrams, instead of (learnable) DeepSet networks, to highlight the benefits provided by the latter. We start by considering a 4-layer GNN, with 3 message passing layers and 1 CliquePH layer, and we change the position of the latter. We consider the structure-based scenario to exclude the influence of node features. Results are shown in Figure 2 (Left) for the DD dataset (other datasets are in the Appendix). We notice that the effects are highly dataset dependent: for DD using CliquePH as the first layer is more effective, while, e.g., for Enzymes the best performance is when the layer is 3rd, similar to PROTEINS. We then consider a GNN with a number of layers that vary from 1 to 5 and with CliquePH as the last layer. Results are shown in Figure 2 (Center). Again we notice that the performance is highly dependent on the dataset, as for DD there is little change in performance for different numbers of layers, while, e.g., for ENZYMES it is important to have 2 to 4 layers prior to CliquePH. On IMDB-B we notice how the addition of CliquePH significantly increases performance even for a 1-layer architecture. Finally, we notice how in all cases the addition of CliquePH always leads to performance improvements. Finally, we apply CliquePH with a maximum dimension of the lifted clique graphs varying from 3 to 5 on the DD dataset. Results are shown in Figure 2 (Right). We observe that tuning this parameter is important to avoid overfitting. These results suggest that tuning the position of CliquePH, and the value of $r$, is important to obtain the best results, and that CliquePH always leads to performance improvements when no node features are available.
@@ 6 Related Work
Graph neural networks and topological data analysis have become very popular and prolific research fields. In this section, we review works at the intersection of graph machine learning and topology, and we refer the interested reader to comprehensive surveys for the latest advancements in graph neural networks [1, 2] and topological data analysis [33, 34, 35].
In the context of graph neural networks augmented with topological information, Bouritsas et al. [7] show that adding information about the number of higher-order structures that a node belongs to can improve the theoretical expressivity and practical performance of a model. Similarly, Immonen et al. [36] use a topological descriptor based on filtrations as additional node features for GNNs. Zhao et al. [37] use persistent homology to weight messages in GNNs in the message passing procedure. He et al. [38] use sheaf theory to produce node position encodings that improve the performance of graph neural networks on several tasks. Song et al. [39] propose a feature augmentation method by obtaining topology embedding of nodes through Node2vec. This introduces topological structure information into an end-to-end model to improve node representation learning.
In addition, several works discuss message-passing on topological spaces. Bodnar et al. [40] introduce a model that performs message passing over simplicial complexes. Bodnar et al. [41] successively extended this to cellular complexes. Later work [42] has also introduced models operating on cellular sheaves. We further mention the work of Hajij et al. [43] which provides a unifying framework for deep learning on topological domains. While these works are related to CliquePH, there are some important differences. Firstly, our method keeps the message passing on the original graph. Secondly, these methods require a specific model, while our method CliquePH can be applied in conjunction with any existing GNN. Thirdly, CliquePH makes use of Persistent Homology, which provides access to information (e.g., the exact number of cliques and cycles in the original graph and in each lifted graph) that is not guaranteed to be accessible to the above methods.
Finally, several works also focus on the explicit connections between topological data analysis and graph learning. O’Bray et al. [44] use graph filtrations to compute efficient graph representations that can outperform those obtained from GNNs. Southern et al. [45] leverage topological tools to robustly evaluate generative graph models. Coupette et al. [46] introduces the concept of curvature for hypergraphs and show that they can be used to effectively perform clustering. Carriere et al. [47] use extended persistent diagrams to produce graph embeddings. Srambical and Rieck [48] use graph filtrations to classify dynamic graphs. Ye et al. [49] and Zhang et al. [50] introduce the use of extended persistent homology, a variant of persistent homology, into graph-related tasks.
@@ 7 Conclusions
Author Contributions
D.B. and B.R. have come up with the initial idea and formalization. D.B. has coordinated the execution of the project. D.B. has implemented the first version of the code. F.S. has contributed to fixing bugs in the code. F.S. has performed the experiments. F.S. has created the plots and tables for the experimental results. D.B. and B.R. have defined the experimental methodology. D.B. and B.R. have written the paper. F.S. contributed reviewing and editing the paper. B.R. has managed the funding acquisition.
Acknowledgements
B.R. was partially supported by the Bavarian state government with funds from the Hightech Agenda Bavaria. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI). F.S acknowledges financial support by The Helmholtz Association of German Research Centres.
@@ Appendix C Full Ablation results
Appendix D Learnable vs. Static Embeddings for Persistence Diagrams
In this section, we investigate the effect of using a (learnable) DeepSet network instead of static embeddings functions for persistence diagrams. In more details, for the static embeddings we consider the following functions: relational hat [53], triangle point transformation, Gaussian point transformation, and line point transformation [47]. We report results in Table 5. We notice that learnable embeddings provide the highest results, and that static embeddings functions can even hurt performance in some datasets.
Table 5: We compare two approaches for embedding persistence diagrams: a (learnable) DeepSet model and a method based on static (non-learnable) embeddings. Following standard practice, we report ROC-AUC scores on the OGBG-molhiv dataset using Graph Convolutional Networks (GCN) for evaluation.
Graph classification Method ENZYMES PROTEINS-FULL IMDB-B MNIST REDDIT-5K OGBG-molhiv GCN-4 ${58.3}\pm{6.1}$ ${76.1}\pm{2.4}$ ${68.6}\pm{4.9}$ ${90.0}\pm{0.3}$ ${53.7}\pm{1.7}$ ${71.9}\pm{1.1}$ GCN-3-CliquePH-1 (Static) ${44.0}\pm{7.0}$ ${77.3}\pm{2.9}$ ${70.4}\pm{4.8}$ ${93.8}\pm{0.5}$ ${50.4}\pm{2.2}$ ${74.5}\pm{1.9}$ GCN-3-CliquePH-1 (Ours; DeepSet) ${55.5}\pm{4.2}$ ${82.0}\pm{0.5}$ ${71.1}\pm{3.2}$ ${95.9}\pm{0.2}$ ${55.6}\pm{1.2}$ ${75.2}\pm{2.3}$
Appendix E Node Classification Experiments
@@
Appendix F Architecture & Datasets Statistics
We report statistics for the datasets in Table 7.
In Tables 9, 9, 10, we report the parameter count for the considered GCN-TOGL, GCN-CliquePH, and base GCN. Notice how the number of parameters remains almost equal between all models. This is due to our choice of replacing a GCN layer with a topological layer (rather than adding a layer on top), and ensures a fairer comparison.
Table 7: Dataset statistics
Graph classification Method #graphs #nodes #edges #features #classes max clique size DD 11781178{1178}1178 284.32284.32{284.32}284.32 1431.321431.32{1431.32}1431.32 8989{89}89 22{2}2 77{7}7 ENZYMES 600600{600}600 32.6332.63{32.63}32.63 124.27124.27{124.27}124.27 33{3}3 66{6}6 55{5}5 MNIST 5500055000{55000}55000 70.5670.56{70.56}70.56 564.50564.50{564.50}564.50 11{1}1 1010{10}10 88{8}8 PROTEINS 11131113{1113}1113 39.0639.06{39.06}39.06 145.63145.63{145.63}145.63 33{3}3 22{2}2 55{5}5 IMDB-B 10001000{1000}1000 19.7719.77{19.77}19.77 193.06193.06{193.06}193.06 00{0} 22{2}2 3030{30}30 REDDIT-5K 49994999{4999}4999 508.52508.52{508.52}508.52 1189.751189.75{1189.75}1189.75 00{0} 55{5}5 66{6}6 OGBG-molhiv 4112741127{41127}41127 25.5125.51{25.51}25.51 54.9454.94{54.94}54.94 99{9}9 22{2}2 44{4}4 Coauthor CS 11{1}1 18333.0018333.00{18333.00}18333.00 163788.00163788.00{163788.00}163788.00 68056805{6805}6805 1515{15}15 2020{20}20 Coauthor Physics 11{1}1 34493.0034493.00{34493.00}34493.00 495924.00495924.00{495924.00}495924.00 84158415{8415}8415 55{5}5 1212{12}12 Cora ML 11{1}1 2995.002995.00{2995.00}2995.00 16316.0016316.00{16316.00}16316.00 28792879{2879}2879 77{7}7 77{7}7
Table 8: GCN-3-CliquePH-1 Model Parameters and Summary
Model Components
Name Type Params
Embedding Linear 1.2 M
Layers ModuleList 125 K
Classif Sequential 13.3 K
CliquePH SimpleSetTopoLayer 34.5 K
Summary
Trainable params 1.4 M
Non-trainable params 0
Total params 1.4 M
Total estimated model params size 5.540 MB
Table 9: GCN-3-TOGL-1 Model Parameters and Summary
Model Components
Name Type Params
Embedding Linear 1.2 M
Layers ModuleList 65.3 K
Classif Sequential 13.6 K
Togl SimpleSetTopoLayer 16.1 K
Summary
Trainable params 1.3 M
Non-trainable params 0
Total params 1.3 M
Total estimated model params size 5.295 MB
Table 10: GCN Model Parameters and Summary
Model Components
Name Type Params
Embedding Linear 1.2 M
Layers ModuleList 87.0 K
Classif Sequential 13.6 K
Summary
Trainable params 1.3 M
Non-trainable params 0
Total params 1.3 M
Total estimated model params size 5.317 MB